OpenStudy (anonymous):

solve... 4x^4-16x^2+15=0

OpenStudy (anonymous):

\[4x^4-10x^2-6x^2+15=0\]

OpenStudy (anonymous):

\[2x^2(2x^2-5)-3(2x^2-5)=0\]

OpenStudy (anonymous):

\[(2x^2-3)(2x^2-5)\]

OpenStudy (anonymous):

x^2=3/2 or x^2=5/2

OpenStudy (anonymous):

x=sqrt(3/2) 0r sqrt(5/2)